OpenStudy (anonymous):

The perimeter of a rectangle 48 is feet. If the length measures 16 feet, which is the width? a-10ft b-5ft c-8ft d-16ft

OpenStudy (anonymous):

why the flutter someone still uses feets in computing? 2a+2b=48 a=16 ---------- 32+2b=48 2b=16 b=8 width is 8 FEET... omg

OpenStudy (anonymous):

well considering that a rectangle perimeter is equal to = 2l + 2w if the l = 16 it be like this 48 = 2(16) + 2(w) 48 = 32 + 2(w) 48 - 32 =2(w) 16 = 2w 16/2 = w 8 = w the width equals 8 feets

OpenStudy (anonymous):

lol I meant why the F.U.C.K and it edited it :D